OpenStudy (anonymous):

Jan is one year younger than her brother Bill and one year older than her sister Sue, The sum of all their ages are 57. How old is each family member?

hero (hero):

Bill = j + 1 Jan = j Sue = j - 1 Bill + Jan + Sue = 57 (j+1) + j + (j-1) = 57 3j = 57 j = 57/3 j = 19 j+1 = 20 j -1 = 18